A 11.4 mL sample of 0.180 M sulfuric acid (H2SO4) is titrated with 0.250 M NaOH Calculate the volume, in milliliters, of NaOH solution that will be used.
Answer #1

Sol.

Balanced reaction :

2NaOH + H2SO4 ---> Na2SO4 + 2H2O  

As Millimoles of H2SO4 = Conc. of H2SO4 × Volume of H2SO4 = 0.180 × 11.4 = 2.052 mmol

From reaction , Millimoles of NaOH = 2 × Millimoles of H2SO4 = 2 × 2.052 = 4.104 mmol

As Conc. of NaOH = 0.250 M

So , Volume of NaOH

= Millimoles of NaOH / Conc. of NaOH

= 4.104 / 0.250

=   16.4 mL